Year 1: Forward Rate is 4.6 %
Year 2: Forward Rate is 4.9 %

Year 3: Forward Rate is 5.2 %

Year 4: Forward Rate is 5.5 %

Year 5: Forward Rate is 6.8 % 


What is the yield to maturity of a 3-year bond?
Answer #1

Given that,

Year 1: Forward Rate f(0,1) = 4.6%

Year 2: Forward Rate f(1,1) = 4.9%

Year 3: Forward Rate f(2,1) = 5.2%

So, Yield to maturity of a 3 year bond is

YTM = ((1+f(0,1))*(1+f(1,1))*(1+f(1,2)))^(1/3) - 1 = (1.046*1.049*1.052)^(1/3) - 1 = 4.90%

So, YTM on 3-year bond is 4.90%